The sum of Bob, Mary and Sue's ages is fifty nine. Bob is twice Sue's age. Mary is one year younger than Sue. Find the age of each person.
Answer by macston(5194)

S=Sue's age; B=bob's age=2S; M=Mary's age=S-1yr;
B+M+S=59yrs Substitute for B and S
2S+S-1yr+S=59yrs Add 1 yr to each side.
4S=60yrs Divide each side by 4
S=15 years ANSWER 1: Sue is 15 years old.
B=2S=2(15yrs)=30 yrs ANSWER 2: Bob is 30 years old.
M=S-1 yr=15 yrds-1yr=14 years ANSWEr 3: Mary is 14 years old.
CHECK:
B+M+S=59yrs
30yrs+14yrs+15yrs=59yrs
59yrs=59yrs